OpenStudy (anonymous):

Ravi runs 6 miles in 45 minutes. At the same rate, how many minutes would he take to run 4 miles?

OpenStudy (anonymous):

\[\frac{ 6 miles }{ 45 minutes }=\frac{ 4 miles }{ ? minutes }\] you just cross multiply to find the missing number

OpenStudy (imstuck):

You could set this up as an equality statement. Like this\[\frac{ 6mi }{ 45\min }=\frac{ 4mi }{ x \min }\]To solve, cross multiply to get 6x = 4(45). 6x = 180. x = 30. So he can run 4 miles in 30 minutes. See that?

OpenStudy (anonymous):

Thank you i think i understand. can you tell me if i am doing this problem correctly. Kira drove 387 miles using 18gallons of gas. At this rate, how many miles would she drive using gallons of gas? I came up with 1.95 gallons

OpenStudy (anonymous):

Is it asking how many miles per gallons shes using?

OpenStudy (anonymous):

if so, you just divide 387 by 18 which equals 21.5 miles per gallon of gas